OpenStudy (anonymous):

Find the zeros write in factored form f(x)=7x^3-x^2+7x-1

OpenStudy (anonymous):

if u put 1/7 in the equation u get 0 means 1/7 is a root so x=1/7 or 7x-1 is a factor now divide the equation by 7x-2 to get other factors

OpenStudy (anonymous):

You have to find the zeros first

OpenStudy (anonymous):

factors first then from that u find zeros

OpenStudy (anonymous):

So divide the whole equation by 7x-2

OpenStudy (whpalmer4):

You can find the zeros by factoring: \[(7x^3 - x^2 ) + (7x - 1)\]Are there any common factors in the left hand parentheses? Factor them out.

OpenStudy (whpalmer4):

After doing that, you should be able to factor again by using the distributive property in reverse: \[ab + ac = a(b + c)\]except here the factor will be a binomial.